# From Instruction to Event: Sound-Triggered Mobile Manipulation

### v1

@@ Abstract
Machine Learning, ICML
Figure 1: Motivation for our work. (Upper) An example of sound-triggered mobile manipulation, the trigger signal is the sound of the event rather than instructions from humans. (Left) Instruction-based mobile manipulation needs humans to analyze the trigger signal and manually give the instruction for the robot model, which is passive. (Right) In sound-triggered mobile manipulation, the robot model automatically receives sound signals and actively outputs executable actions.
1 Introduction
Recent advances in Embodied AI have witnessed remarkable progress in mobile manipulation, enabling robots to navigate unstructured environments and interact with physical objects (Song et al., 2025a; Bai et al., 2025; Yu et al., 2025). Such synergy between mobility and interaction empowers agents to perform long-horizon household tasks such as object rearrangement (Ji et al., 2025; Huang et al., 2025; Escoriza et al., 2025) and multi-stage navigation (Song et al., 2025b; Lin et al., 2025), bringing us closer to general-purpose home robotic assistants.
Current paradigm predominantly studies mobile manipulation under an instruction-conditioned setting, where the task goal and intended interaction are explicitly specified by a natural language instruction (Wu et al., 2025; Shah et al., 2025; Yan et al., 2025; Zhang et al., 2025). While effective for user-specified chores, this paradigm is ill-suited for the stochastic nature of daily life. For example, events such as a boiling kettle or a ringing doorbell typically happen unexpectedly and require timely intervention, where the latency inherent in human-in-the-loop interaction renders timely intervention impossible. This uncertainty is further compounded when multiple events occur concurrently or in rapid succession, i.e., waiting for a phone call from friends while expecting a doorstep delivery from a courier simultaneously. The indeterminate sequence of such events prevents the user from formulating a valid instruction plan in advance.
This limitation underscores the discrepancy between static, predefined instructions and dynamic, sound-triggered scenarios (See Figure 1). Although recent efforts have attempted to generalize instructions ranging from concrete commands (Yenamandra et al., 2023) to abstract goals (Wu et al., 2023) and ambiguous queries (Ichter et al., 2022; Li et al., 2025), addressing these stochastic events remains an open challenge. In this context, we find that sound signals play a pivotal role. For instance, if the phone rings prior to a knock at the door, the robot should prioritize picking up the phone before retrieving the delivery. In such cases, robotic actions are modulated by stochastic sounds rather than predefined text instructions. Specifically, sounds serve a threefold function: (1) temporal trigger indicating the start of a stochastic event; (2) semantic cue guiding the robot’s action selection; and (3) spatial cue revealing the event’s location. Based on these insights, we introduce sound-triggered mobile manipulation, comprising SonicStow, SonicInteract, and Bi-Sonic Manipulation, where acoustic signals replace conventional textual instructions as the trigger signal. SonicStow and SonicInteract present unique challenges beyond traditional visual manipulation: (1) Acoustic semantic understanding, which requires the robot to interpret sounds to determine target locations and actions. (2) Sound-conditioned interaction, which demands the execution of articulated interactions solely triggered by audio. The Bi-Sonic Manipulation task further elevates the complexity by introducing dual sound sources, which demands the agent to prioritize the primary source amidst acoustic interference and subsequently locate and interact with the secondary sound source.
To validate our proposed tasks, we introduce Habitat-Echo, a simulation platform that seamlessly integrates acoustic rendering with interactive physical environments. Built upon the visual and physical capabilities of Habitat (Szot et al., 2021; Puig et al., 2024), Habitat-Echo addresses the absence of sound signals in prior simulators. Specifically, we leverage room impulse responses to convolve source sound with the environment’s geometry, thereby ensuring physically accurate sound propagation and spatialization from the source to the receiver. Leveraging this platform, we construct a benchmark tailored for sound-triggered mobile manipulation. The dataset encompasses many interactive assets, categorized into: (1) static silent rigid objects, (2) sound-emitting rigid objects, and (3) sound-emitting articulated objects. We utilize this dataset to train policy models and further introduce an Omni-modal Large Language Model (Omni-LLM) as a high-level task planner to understand acoustic events and decompose complex manipulation goals. The main contributions of this work are summarized as follows:
  • Novel Sound-Triggered Tasks. Distinct from traditional instruction-following paradigms, we formulate three novel manipulation tasks triggered by stochastic acoustic signals. In these tasks, robotic actions are autonomously triggered by stochastic sound events rather than static predefined instructions, better reflecting the unpredictable and dynamic nature of real-world household scenarios.
  • New Audio-Visual-Physical Simulation Platform. We introduce Habitat-Echo, a platform that integrates acoustic and visual rendering with interactive physical environments. By incorporating sound-emitting objects, Habitat-Echo facilitates large-scale episode generation and closed-loop online interaction, establishing a robust foundation for future research in embodied audio-visual learning.
  • Hierarchical Baseline and Validation. We develop a hierarchical framework that is composed of a high-level task planner to interpret acoustic cues and several low-level executors for precise control. We leverage an Omni-LLM as the task planner and policy models as executors. The baseline framework achieves promising performance on the proposed benchmark.
@@ 1 Introduction
Figure 2: Illustration of the proposed tasks. (a) SonicStow requires the agent to interact with the sound source (a rigid object) via Navigate, Pick, and Place skills. (b) SonicInteract requires the agent to interact with the sound source (an articulated object) via Navigate, Open Door, and Close Sink skills. (c) Bi-Sonic Manipulation requires the agent to interact with the sound source (two objects sequentially) via Navigate, Pick, Place, Open Door, and Close Sink skills.
2 Related Work
Sound Navigation. Sound Navigation, proposed by (Chen et al., 2020), requires an agent to navigate to an audio source without explicit goal coordinates. The primary challenges involve understanding physical spaces, localizing sound emitters, and avoiding obstacles. Existing methods can be broadly categorized into two streams: real-world application and versatile training. For real-world application, researchers aim to bridge the simulation-to-reality gap by introducing complex environmental settings. (Chen et al., 2021) assign semantic meanings to audio sources and introduce dynamic sound emission, replacing the continuous, generic signals used in prior work. (Younes et al., 2023) further extends the setting from static to moving sound sources. To simulate complex acoustic environments, (Yu et al., 2022) introduces an adversarial sound attacker that generates distracting noises with varying categories and volumes. For versatile training, methods leverage cross-modal synergies to enhance performance. VXN (Wang et al., 2022) unifies diverse navigation tasks, using sound navigation to facilitate the training of visual-language navigation. Recently, (Kondoh and Kanezaki, 2025) incorporates language descriptions of past and future actions as an auxiliary task to enhance policy learning. However, these works mainly focus on navigation capabilities, neglecting physical interactions with objects within the environment.
@@ 2 Related Work
Task Planning for Robotic Manipulation. Task planning bridges high-level reasoning with low-level execution. Early approaches (Kaelbling and Lozano-Pérez, 2011; Toussaint, 2015; Srivastava et al., 2014) primarily relied on predefined symbolic representations for continuous control. Recent research leverages the extensive world knowledge within LLMs and VLMs for planning, broadly categorized into two streams The first stream focuses on task decomposition. Wake et al. (Wake et al., 2024) utilize GPT-4V (OpenAI, 2023) to ground spatial-temporal information within the demonstration video and analyze affordances for plan generation. Similarly, Kachaev et al. (Kachaev et al., 2025) employ VLMs to decompose instructions into subtask sequences. RoboBrain (Ji et al., 2025) unifies planning with affordance perception and trajectory prediction within a single model, enhancing manipulation capabilities. The second stream emphasizes feedback-based refinement. Zhi et al. (Zhi et al., 2025) optimize code-based plans using execution feedback, and Feng et al. (Feng et al., 2025) improve plans by anticipating future states. Similarly, Wang et al. (Wang et al., 2025) employ VLMs to solve PDDL problems constructed from current observations. However, these works primarily rely on visual and textual inputs, neglecting the importance of audio signals.
3 Task Definition and Skill
3.1 Task Definition
We formally define three novel sound-triggered tasks: SonicStow, SonicInteract, and Bi-Sonic Manipulation. SonicStow and SonicInteract are mobile manipulation tasks with a single sound source. Bi-Sonic Manipulation, a scenario with dual sound sources. See Figure 2 for more details.
SonicStow. This task simulates a scenario where an agent actively perceives an acoustic signal, localizes its source, and performs Pick and Place operations on household objects. Distinct from standard instruction-following paradigms that rely on predefined instructions, the agent must autonomously identify the target and conduct interaction solely based on the stochastic audio event. In each episode, the agent is initialized with a random pose and start location, while a single sound source, randomly selected from three objects (Phone, Alarm, Furby) within the scene (see Figure 2(a)). To introduce visual clutter, two distractor objects are randomly sampled from the YCB dataset for each episode. All objects are placed on elevated receptacles (e.g., existing furniture or inserted tables) instead of floor.
SonicInteract. This task entails localizing an acoustic event and performing manipulation on articulated objects. Different from picking and placing objects in SonicStow, this task requires the agent to precisely control the end-effector and manipulate articulated mechanisms, i.e., opening a door or turning off a faucet in the sink (see Figure 2(b)). In each episode, the agent is initialized with a random pose and start location. The target single-sound source is randomly sampled from two categories: a doorbell or a sink with running water. Object placement follows category-specific constraints: doorbells are fixed to the existing doors within the scene, whereas sinks are placed on the floor with arbitrary orientations and locations.
Bi-Sonic Manipulation. The Bi-Sonic Manipulation task integrates the challenges of SonicStow and SonicInteract, requiring the agent to manipulate both rigid and articulated objects (see Figure 2(c)). Additionally, this task introduces three more challenges. The first challenge of this task is the presence of two simultaneous sound sources, which creates a complex acoustic field featuring signal dynamics and superposition. Unlike single-source tasks, the agent must distinguish and interact with both sources sequentially based on a given priority (i.e., the main source followed by the secondary source). The categories and spatial locations of these sources are randomized. The second challenge of this task is two-time interactions. After interacting with the first object, the sound of the first object stops, and there is only one sound source in the scene. The agent has to execute skills and interact with the second object. Furthermore, to enhance visual complexity, we still introduce unrelated objects (Calli et al., 2015) to serve as visual clutter.
3.2 Skill
Similar to previous work (Gu et al., 2023; Shukla et al., 2024; Wang et al., 2025), the proposed tasks are decomposed into different subtasks. Each subtask is solved by an individual skill. A task is considered completed successfully if and only if all constituent skills are successfully executed in sequence. There are five skills in the proposed three tasks.
Navigate. The action space consists of four actions: MoveForward, TurnLeft, TurnRight, and Stop. Specifically, MoveForward translates the agent by 0.5m along its local forward axis, while rotational actions (TurnLeft, TurnRight) execute $90^{\circ}$ in-place rotations to the left and right, respectively. The Stop action terminates the episode without changing the agent’s pose and location. An episode is considered successful if and only if the agent executes Stop while positioned within a region of 0.5m radius of the goal.
Pick. The agent operates in an 8-DoF continuous action space. The first seven dimensions control the arm’s joint angles, driving the end-effector to the target pose via the kinematic chain (Abdelmaksoud et al., 2024), while the final dimension modulates the gripper state. A Pick skill is considered successful with two conditions: (1) Grasping: The gripper action is positive while an object is within a 15cm snapping threshold of the end-effector, triggering a stable grasp. (2) Retraction: The agent effectively holds the object and returns the arm to a predefined resting pose, ensuring a smooth transition to subsequent subtasks.
Place. The agent operates in an 8-DoF action space identical to that of the Pick skill. The success conditions for the Place skill are twofold: (1) Releasement: The gripper opens to release the object, ensuring it settles within 0.15cm of the target position. (2) Retraction: Upon release, the arm reverts to the default resting pose, preparing for the next subtask.
Open Door. Operating within the 8-DoF action space, the agent’s end-effector attaches to the handle of the door and drives it along its own functional trajectory. The skill is considered successful with two conditions: (1) Opening: While connected with the door handle, the agent must drive the door’s joint angle to a specified target threshold. (2) Retraction: The agent disengages from the door handle and returns the arm to the resting pose.
Close Sink. Similar to the Open Door skill, the agent utilizes the 8-DoF action space to manipulate the sink’s rotary mechanism. The end-effector must grasp the faucet handle and apply torque to rotate it along its functional axis. Success is determined by two conditions: (1) Closure: While maintaining grasp, the agent must rotate the handle to the zero-degree threshold within a tolerance of 0.2 rad. (2) Retraction: Upon completing the rotation, the agent releases the handle and reverts the arm to the resting pose.

### v2

@@ Abstract.
Audio, Mobile Manipulation, Task Planning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 10–14, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06submissionid: 7623ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Robotic planning
Figure 1. Motivation for our work. This scenario presents two simultaneous and stochastic events triggered by sounds: a ringing doorbell and an alarm. The agent is required to actively explore the environment. Unlike the conventional instruction-driven paradigm, Sound-Triggered Mobile Manipulation (STMM) can effectively handle unpredictable events without relying on explicit textual instructions.
1. Introduction
Recent advances in embodied AI have led to remarkable progress in mobile manipulation, enabling robots to navigate in unstructured environments and physically interact with surrounding objects (Song et al., 2025a; Bai et al., 2025; Yu et al., 2025). Such synergy between mobility and interaction empowers agents to perform long-horizon household tasks such as object rearrangement, multi-stage interaction, and sequential task execution (Ji et al., 2025; Huang et al., 2025; Li et al., 2025b; Escoriza et al., 2025; Song et al., 2025b; Lin et al., 2025), pushing embodied agents closer to general-purpose household assistance. Despite these advances, most existing mobile manipulation works are built upon an instruction-driven paradigm, where the task goal is explicitly specified by the user through instructions (Wu et al., 2025; Shah et al., 2025; Yan et al., 2025; Zhang et al., 2025; Liu et al., 2025a; Li et al., 2025a), and the agent is expected to execute the instructed behavior accordingly. While this paradigm excels at completing user-specified chores, it also implicitly assumes that the task objective is known before execution begins. However, in real-world household environments, many tasks are triggered by stochastic events that require a timely response. For example, a ringing doorbell may prompt the agent to approach the door and open it, and a ringing alarm may indicate that the device needs to be turned off (see Figure 1). The unpredictable nature of these events makes it hard for users to monitor the situation and provide explicit instructions in time. This highlights the urgent need for event-driven mobile manipulation, where an intelligent agent should be able to proactively perceive and respond to stochastic environmental events.
Among various environmental signals, sound serves as a natural trigger for event-driven interaction. Sound is omnidirectional and independent of illumination compared to visual cues. In fact, many of these events are exposed acoustically long before they enter the visual field. An agent can therefore detect critical events even when the sounding target lies outside its field of view or under poor lighting conditions. Moreover, sound signals not only indicate the occurrence of an event, but also provide partial semantic cues about the event type and spatial cues about the location of the sound source. These properties make sound a uniquely informative modality for proactive embodied response. Motivated by this observation, we introduce a novel problem termed Sound-Triggered Mobile Manipulation (STMM), where the agent must autonomously ground the sound source and execute appropriate interactions based on the auditory event and scene context. Unlike conventional instruction-driven mobile manipulation, STMM requires the agent to infer the task goal from environmental events rather than receiving it directly from a human user (see Figure 1).
Concretely, we categorize sound-triggered mobile manipulation into two fundamental interaction types: (1) object relocation, where the agent localizes a sounding rigid object (e.g., a phone) and alters its spatial position via pick-and-place skills, and (2) state transition, where the agent localizes a sounding functional or articulated object (e.g., a door) and manipulates it to change its internal physical states, such as opening or closing. Building upon this, we investigate both single-event and more challenging multi-event settings. In the multi-event setting where multiple sound events occur simultaneously, the agent must not only determine the correct interaction for each event, but also resolve event prioritization and execute tasks in an appropriate order. Overall, sound-triggered mobile manipulation demands the agent to jointly reason about what event is happening, where the relevant sounding target is located, and how the correct interaction should be carried out, while additionally handling task scheduling in multi-event scenarios, making the problem particularly challenging.
However, existing embodied simulators either focus strictly on visual-physical interaction (e.g., mobile manipulation simulator (Puig et al., 2024)) or visual-sound sensing (e.g., sound navigation environment (Chen et al., 2020)). There is a lack of an integrated environment where agents can perceive spatial sound and physically interact with sound-emitting targets within a closed loop, thus significantly hinders systematic research on sound-triggered mobile manipulation. To bridge this gap, we develop Habitat-Echo, a unified audio-visual-physical simulation platform. By seamlessly integrating spatial audio rendering into existing interactive embodied environments and extending it with abundant sound-emitting interactive assets, Habitat-Echo enables a closed-loop simulation of sound perception, sound source grounding, and physical interaction within a single unified platform.
Built on Habitat-Echo, we further establish STMM-1.9K, the first benchmark for sound-triggered mobile manipulation, covering diverse sounding objects and interaction scenarios under both single-event and multi-event settings. To provide a first-step solution to this new problem, we design a hierarchical baseline framework that decomposes sound-triggered mobile manipulation into high-level acoustic reasoning and low-level physical execution. For high-level acoustic reasoning, a task planner powered by an Omni-MLLM (Xu et al., 2025) is leveraged to translate audio-visual observations into an executable skill chain, including task priority and the sequence of required skills. For low-level physical execution, a set of lightweight policy models is invoked sequentially to carry out each required skill. Each policy model focuses on a specific skill, such as navigation, pick-and-place, and door-opening, enabling specialized control for diverse interaction types. Experimental results on the STMM-1.9K benchmark provide an initial validation of the proposed setting and reveal its core challenges, including auditory grounding, task prioritization, and multi-stage execution.
Our contributions are summarized as follows:
  • We propose Sound-Triggered Mobile Manipulation (STMM), a novel event-driven embodied task in which agents are required to respond to sound events physically without relying on explicit per-episode instructions.
  • We develop Habitat-Echo, a unified embodied simulation platform that seamlessly unifies high-fidelity spatial audio perception with interactive physical interaction.
  • We establish STMM-1.9K, the first benchmark for STMM, and propose a hierarchical baseline framework that offers an initial solution and reveals the core challenges of this new embodied task.
@@ 1. Introduction
Figure 2. Different settings of Sound-Triggered Mobile Manipulation (STMM). (a) Object Relocation in the single-event setting requires the agent to interact with the sound source (a rigid object) via Navigate, Pick, and Place skills. (b) State Transition in the single-event setting requires the agent to interact with the sound source (an articulated object) via Navigate, Open Door, and Close Sink skills. (c) Multi-event Setting requires the agent to interact with the sound sources sequentially.
2. Related Work
@@
Sound in Embodied AI. Sound provides crucial spatial and contextual information for embodied agents. Existing applications of sound in embodied AI can be broadly categorized into two main areas: navigation and manipulation. Audio-Visual Navigation, pioneered by (Chen et al., 2020), requires an agent to navigate to an audio source without explicit goal coordinates. Existing methods generally follow two streams: complex environmental modeling for real-world applications and cross-modal strategies for versatile training. For sim-to-real gap, (Chen et al., 2021) replaces continuous, generic signals with dynamic, semantically meaningful sound emissions (Chen et al., 2021). This has been further extended to moving sound sources (Younes et al., 2023) and complex acoustic environments featuring adversarial attackers that generate distracting noises (Yu et al., 2022). For versatile training, VXN (Wang et al., 2022) unifies diverse navigation tasks by using sound navigation to facilitate vision-language training. (Kondoh and Kanezaki, 2025) incorporates language descriptions of past and future actions as an auxiliary task to enhance policy learning. However, these works predominantly focus on navigation capabilities, neglecting fine-grained physical interactions with objects. In robotic manipulation, sound typically serves as an informative modality to guide physical processes. (Wang et al., 2025a) utilizes MMAudio (Ruan et al., 2023) to generate the sound of pouring liquid into a bottle to assist the process. RoboOmni (Wang et al., 2026) extends textual instructions into audio contexts, enabling interactive communication with humans. Additionally, (Guo et al., 2026) employs beamforming to convert acoustic signals into 2D masks, complementing visual observations. Nevertheless, these approaches primarily treat sound as a complementary input, rather than the primary triggering signal that initiates an event.
@@
Task Planning. Recent LLM advancements greatly broaden the scope of task planning. Specifically, Tool-Planner (Liu et al., 2025b) minimizes error propagation in tool usage by effectively grouping and selecting APIs. In algorithmic planning, (Katz et al., 2024) boosts planning efficiency while maintaining logical rigor through optimized tree-based methods. For web agents, (Wang et al., 2025b) proposes a benchmark evaluating LLMs’ ability to translate user intentions into step-by-step product associations. Task planning in robotic manipulation bridges high-level reasoning with low-level execution. Early approaches (Kaelbling and Lozano-Pérez, 2011; Toussaint, 2015; Srivastava et al., 2014) primarily relied on predefined symbolic representations for continuous control. Recent research leverages the extensive world knowledge within LLMs and VLMs for planning, broadly categorized into two streams The first stream focuses on task decomposition. Wake et al(Wake et al., 2024) utilize GPT-4V (OpenAI, 2023) to ground spatial-temporal information within the demonstration video and analyze affordances for plan generation. RoboBrain (Ji et al., 2025) unifies planning with affordance perception and trajectory prediction within a single model, enhancing manipulation capabilities. The second stream emphasizes feedback-based refinement. Zhi et al(Zhi et al., 2025) optimize code-based plans using execution feedback, and Feng et al(Feng et al., 2025) improve plans by anticipating future states. However, these works primarily rely on visual and textual inputs, neglecting the importance of audio signals.
3. Sound-Triggered Mobile Manipulation
3.1. Problem Formulation
Unlike conventional instruction-driven mobile manipulation, which is typically formulated as learning a policy $\pi(a_{t+1}\mid O_{t},L)$ conditioned on an explicit instruction $L$, Sound-Triggered Mobile Manipulation (STMM) removes the explicit instruction that specifies the task goal. Instead, we formulate STMM as a partially observable Markov decision process. At each time step $t$, the agent only receives a multimodal observation $O_{t}=\{V_{t},A_{t},P_{t}\}$, where $V_{t}$, $A_{t}$, and $P_{t}$ denote visual observations, sound observations, and proprioceptive states, respectively. Triggered by environmental sound events $E_{sound}$, the agent must implicitly infer the task goal $\mathcal{G}$ from its observations and execute actions $a_{t}\in\mathcal{A}$ over time to interact with the relevant target. The episode is considered successful if the event-conditioned goal induced by $E_{\text{sound}}$ is successfully accomplished.
Concretely, as shown in Figure 2, depending on the physical properties of the sound-emitting targets associated with $E_{sound}$, we categorize the required interactions to fulfill $\mathcal{G}$ into two fundamental types:
Object Relocation. Here, the sounding target is a rigid object (e.g., a ringing phone or alarm in Figure 2(a)), and the goal is to resolve the event by changing its spatial position (e.g., retrieving the ringing phone). To accomplish this, agent must localize and approach the relevant sound source, and then perform object-centered manipulation such as grasping and placing. The main challenge lies in translating cross-modal grounding to executable object interaction.
State Transition. Here, the sounding target is a functional or articulated object governed by object-specific kinematic constraints (e.g., a running faucet or a door associated with a ringing doorbell in Figure 2(b)), and the goal is to resolve the event by changing its internal physical state (e.g., turning of the faucet). Compared to object relocation, since the agent must operate along the object’s functional axis or articulated trajectory, this interaction type imposes stronger constraints on local geometry and manipulation precision. For example, the agent must precisely grasp a faucet handle and rotate it along its functional axis to stop the running water.
3.2. Task Setting
Beyond the interaction types defined above, we further design single-event and multi-event settings for STMM, which differ in the complexity of source localization and decision making.
Single-event Setting. In the single-event setting, only one sounding event is present in the scene. The agent is required to infer the event-conditioned goal from the multimodal observation, localize the relevant target, and execute the corresponding interaction to resolve the event. Depending on the physical properties of the sounding target, the required interaction can take the form of either object relocation or state transition.
Multi-event Setting. In the multi-event setting, multiple sound events occur simultaneously in the same scene, leading to a more complex acoustic field with signal superposition and source interference. Compared with the single-event setting, the agent must not only isolate the primary sound source from interference but also perform priority reasoning to execute the required interactions in an appropriate order. After the first event is resolved, its corresponding sound disappears and the acoustic scene changes, requiring the agent to continue interacting with the remaining target under an updated environment state.
3.3. Skill Space
In STMM, directly learning a policy from raw multimodal observations to low-level actions inextricably entangles high-level event reasoning with low-level physical execution, which imposes severe learning burdens and hinders proactive decision-making. To mitigate this, we decouple these complex embodied responses into a set of reusable skill primitives, where each skill represents an atomic sensorimotor mapping tailored to a specific physical interaction. Specifically, the skill space is define as:
(1) $\mathcal{K}=\{\texttt{Navigate},\texttt{Pick},\texttt{Place},\texttt{OpenDoor},\texttt{CloseSink}\}.$
Here, Navigate denotes approaching the relevant sounding target. Pick and Place correspond to rigid-object relocation via the agent’s gripper. OpenDoor and CloseSink correspond to state-transition interactions. The detailed execution spaces and termination criteria are elaborated in the Supplementary materials. Under this abstraction, different events can be solved through different skill compositions, i.e., the Skill Chain. An event is considered resolved successfully if and only if all constituent skills are successfully executed in sequence. Object relocation is typically instantiated as:
(2) $\texttt{Navigate}\rightarrow\texttt{Pick}\rightarrow\texttt{Place}],$
whereas state transition is instantiated as:
(3) $\texttt{Navigate}\rightarrow\texttt{OpenDoor}]~~or~~[\texttt{Navigate}\rightarrow\texttt{CloseSink}].$
For multi-event scenes, the complete task is further represented as the ordered composition of multiple skill chains with an explicit priority structure. In this way, the skill chain serves as a structured bridge from latent event interpretation to executable embodied responses.
4. Platform and Benchmark
4.1. Habitat-Echo Simulation Platform